OpenStudy (anonymous):

What is the slope of the line that passes through the points (-2-5) and (1,4) A. -3 B. -2 C. -1/3 D. 1/3

OpenStudy (anonymous):

fan and medal :)

OpenStudy (astrophysics):

You can use the slope formula \[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] where \[(x_1,y_1)~~~~(x_2,y_2)\] are your points

OpenStudy (anonymous):

I'm sorry I don't understand please explain ;/

OpenStudy (astrophysics):

It's the slope formula, rise/run

OpenStudy (astrophysics):

m is your slope just plug and chug your points, refer to this http://www.mathsisfun.com/geometry/slope.html

OpenStudy (anonymous):

4-(-5)÷1-(-2)=9/3=3/1= 3, slope = 3

OpenStudy (anonymous):

sorry the slope is -3